# Boundary four-point connectivities of conformal loop ensembles

### 30 Mar 2026 v1

@@ Appendix A Boundary Green’s functions of SLE
In this appendix we show the equivalence of various definitions of SLE boundary Green’s functions. For $\kappa\in(4,8)$, let $\eta$ be a $\operatorname{SLE}_{\kappa}$ on $\mathbbm{H}$ from 0 to $\infty$. The boundary Green’s function in [FZ23] is defined to be the normalized limit of the probability that $\eta$ approaches the neighborhood of $x_{1},...,x_{n}\in\mathbbm{R}$.
@@
where $I(x,r):=[x-r,x+r]\subset\mathbbm{R}$. Furthermore, by [Zha22, Theorem 6.17], $\int_{S}\widetilde{G}(x)dx=\mathbbm{E}\left[\nu_{\eta\cap\mathbbm{R}}(S)\right]$ and $\int_{S_{1}\times S_{2}}\widetilde{G}(x_{1},x_{2})dx_{1}dx_{2}=\mathbbm{E}\left[\prod_{i=1}^{2}\nu_{\eta\cap\mathbbm{R}}(S_{i})\right]$ for any open $S,S_{1},S_{2}\subset\mathbbm{R}$. Hence, $G(x)=\widetilde{G}(x)$ and $G(x_{1},x_{2})=\widetilde{G}(x_{1},x_{2})$ (recall (2.12)). As a corollary of Proposition A.2, we have
@@
Suppose $I\subset\mathbbm{R}$ is an interval, and $\Gamma$ is a $\operatorname{CLE}_{\kappa}$ configuration on $\mathbbm{H}$. For $x\in\mathbbm{R}$, Let $T_{\varepsilon}(x)$ (resp. $\widehat{T}_{\varepsilon}(x)$) be the event that there exists a loop $\ell$ intersecting $I$ and $B(x,\varepsilon)$ (resp. $I(x,\varepsilon)$; in case there are two or more such loops, we e.g. take the leftmost one on $I$). Then the law of ${\bf 1}_{T_{\varepsilon}(x)}\ell$ and ${\bf 1}_{\widehat{T}_{\varepsilon}(x)}\ell$, times $\varepsilon^{-h}$, converges to $C\mu_{x}^{\mathrm{Bub}}$ restricted to intersect $I$ for some constant $C\in(0,\infty)$.
@@
We focus on $\widehat{T}_{\varepsilon}(x)$, and the case for $T_{\varepsilon}(x)$ is similar (with using Lemma A.3). Consider the CLE exploration process by discovering loops that intersect $B(0,\varepsilon)$, then conformally map the remaining unbounded component to $\mathbbm{H}$ (with $I$ fixed), until there exists a loop intersecting both $B(0,\varepsilon)$ and $I$. By the same argument as [SW12, Section 4], as $\varepsilon\to 0$, $\mathbbm{P}[\widehat{T}_{\varepsilon}(x)]$ is $\varepsilon^{h+o(1)}$; furthermore, conditioned on $\widehat{T}_{\varepsilon}(x)$, the conditional law of $\ell$ converges to $\mu_{0}^{\mathrm{Bub}}$ conditioned to intersect $I$. On the other hand, on the event $\widehat{T}_{\varepsilon}(x)$, let $Q_{x,\varepsilon}$ be the law of the leftmost loop on $I$ that intersects with $I(x,\varepsilon)$. Then for any fixed bounded interval $U\subset\mathbbm{R}$ such that $U\cap I=\emptyset$, consider the measure $\mathbbm{N}_{\varepsilon}:=\mathbbm{P}[\widehat{T}_{\varepsilon}(x)]^{-1}{\bf 1}_{x\in U}Q_{x,\varepsilon}(d\gamma)dx$ in the place of $\mathbbm{M}_{\varepsilon}$ in the proof of Proposition 2.2. Then similar to that proof, for compactly supported and continuous $f$, $\mathbbm{N}_{\varepsilon}[f]$ on the one hand converges to $C\int_{U}f(x,\gamma){\bf 1}_{\gamma\cap I\neq\emptyset}\mu^{\mathrm{Bub}}_{x}(d\gamma)dx$ for some $C\in(0,\infty)$, while by the existence of boundary Minkowski content in [Zha22, Theorem 6.17], $\mathbbm{N}_{\varepsilon}[f]$ on the other hand converges to $\left(\lim_{\varepsilon\to 0}\frac{\varepsilon^{h}}{\mathbbm{P}[\widehat{T}_{\varepsilon}(x)]}\right)\cdot\mathbbm{E}\left[\int_{U}\sum_{\ell\in\Gamma}f(x,\ell){\bf 1}_{\ell\cap I\neq\emptyset}\nu_{\ell\cap\mathbbm{R}}(dx)\right]$. In particular, the limit $\lim_{\varepsilon\to 0}\frac{\varepsilon^{h}}{\mathbbm{P}[\widehat{T}_{\varepsilon}(x)]}\in(0,\infty)$ exists. The result then follows. ∎
@@ Appendix B Discrete convergence
The existence of these two limits was shown in [CF24]. For any disjoint $A,B\subset\mathbbm{H}$, we write $\{A\leftrightarrow B\}$ for the event that there exists an open path connecting $A$ and $B$. Without loss of generality, we suppose $\min|x_{i}-x_{j}|>1$. For $s>0$, let $r_{i}<s$. Define
@@
for some constant $C^{\prime},C^{\prime\prime}\in(0,\infty)$. Since $\lim_{\delta\to 0}\frac{\mathbbm{P}^{\delta}[x_{1}^{\delta}\leftrightarrow x_{2}^{\delta}\leftrightarrow x_{3}^{\delta}\leftrightarrow x_{4}^{\delta}]}{\mathbbm{P}^{\delta}[x_{1}^{\delta}\leftrightarrow x_{2}^{\delta},x_{3}^{\delta}\leftrightarrow x_{4}^{\delta}]}$ tends to $1$ as $x_{2}\to x_{3}$, we find $C^{\prime\prime}=C$. Combined with (B.1), we obtain (B.2). ∎

### 06 Apr 2026 v2

@@ Appendix A Boundary Green’s functions of SLE
In this appendix we show the equivalence of various definitions of SLE boundary Green’s functions. For $\kappa\in(4,8)$, let $\eta$ be a $\operatorname{SLE}_{\kappa}$ on $\mathbbm{H}$ from 0 to $\infty$. The boundary Green’s function in [FZ23] is defined to be the normalized limit of the probability that $\eta$ approaches the neighborhood of $x_{1},...,x_{n}\in\mathbbm{R}$. For $n=1,2$, the existence of such boundary Green’s functions was earlier proved in [Law15].
@@
where $I(x,r):=(x-r,x+r)\subset\mathbbm{R}$. Furthermore, by [Zha22, Theorem 6.17], $\int_{S}\widetilde{G}(x)dx=\mathbbm{E}\left[\nu_{\eta\cap\mathbbm{R}}(S)\right]$ and $\int_{S_{1}\times S_{2}}\widetilde{G}(x_{1},x_{2})dx_{1}dx_{2}=\mathbbm{E}\left[\prod_{i=1}^{2}\nu_{\eta\cap\mathbbm{R}}(S_{i})\right]$ for any open $S,S_{1},S_{2}\subset\mathbbm{R}$. Hence, $G(x)=\widetilde{G}(x)$ and $G(x_{1},x_{2})=\widetilde{G}(x_{1},x_{2})$ (recall (2.12)). As a corollary of Proposition A.2, we have
@@
Suppose $I\subset\mathbbm{R}$ is an interval, and $\Gamma$ is a $\operatorname{CLE}_{\kappa}$ configuration on $\mathbbm{H}$. For $x\in\mathbbm{R}$, Let $T_{\varepsilon}(x)$ (resp. $\widehat{T}_{\varepsilon}(x)$) be the event that there exists a loop $\ell$ intersecting II and $B(x,\varepsilon)$ (resp. $I(x,\varepsilon)$; in case there are two or more such loops, we e.g. take the leftmost one on II). Then the law of ${\bf 1}_{T_{\varepsilon}(x)}\ell$ and ${\bf 1}_{\widehat{T}_{\varepsilon}(x)}\ell$, times $\varepsilon^{-h}$, converges to $C\mu_{x}^{\mathrm{Bub}}$ restricted to intersect II for some constant $C\in(0,\infty)$.
@@
We focus on $\widehat{T}_{\varepsilon}(x)$, and the case for $T_{\varepsilon}(x)$ is similar (with using Lemma A.3). Consider the CLE exploration process by discovering loops that intersect $B(0,\varepsilon)$, then conformally map the remaining unbounded component to $\mathbbm{H}$ (with II fixed), until there exists a loop intersecting both $B(0,\varepsilon)$ and II. By the same argument as [SW12, Section 4], as $\varepsilon\to 0$, $\mathbbm{P}[\widehat{T}_{\varepsilon}(x)]$ is $\varepsilon^{h+o(1)}$; furthermore, conditioned on $\widehat{T}_{\varepsilon}(x)$, the conditional law of $\ell$ converges to $\mu_{0}^{\mathrm{Bub}}$ conditioned to intersect II. On the other hand, on the event $\widehat{T}_{\varepsilon}(x)$, let $Q_{x,\varepsilon}$ be the law of the leftmost loop on II that intersects with $I(x,\varepsilon)$. Then for any fixed bounded interval $U\subset\mathbbm{R}$ such that $U\cap I=\emptyset$, consider the measure $\mathbbm{N}_{\varepsilon}:=\mathbbm{P}[\widehat{T}_{\varepsilon}(x)]^{-1}{\bf 1}_{x\in U}Q_{x,\varepsilon}(d\gamma)dx$ in the place of $\mathbbm{M}_{\varepsilon}$ in the proof of Proposition $2.$ Then similar to that proof, for compactly supported and continuous $f$, $\mathbbm{N}_{\varepsilon}[f]$ on the one hand converges to $C\int_{U}f(x,\gamma){\bf 1}_{\gamma\cap I\neq\emptyset}\mu^{\mathrm{Bub}}_{x}(d\gamma)dx$ for some $C\in(0,\infty)$, while by the existence of boundary Minkowski content in [Zha22, Theorem 6.17], $\mathbbm{N}_{\varepsilon}[f]$ on the other hand converges to $\left(\lim_{\varepsilon\to 0}\frac{\varepsilon^{h}}{\mathbbm{P}[\widehat{T}_{\varepsilon}(x)]}\right)\cdot\mathbbm{E}\left[\int_{U}\sum_{\ell\in\Gamma}f(x,\ell){\bf 1}_{\ell\cap I\neq\emptyset}\nu_{\ell\cap\mathbbm{R}}(dx)\right]$. In particular, the limit $\lim_{\varepsilon\to 0}\frac{\varepsilon^{h}}{\mathbbm{P}[\widehat{T}_{\varepsilon}(x)]}\in(0,\infty)$ exists. The result then follows. ∎
@@ Appendix B Discrete convergence
The existence of these two limits was shown in [CF24]. For any disjoint $A,B\subset\mathbbm{H}$, we write $\{A\leftrightarrow B\}$ for the event that there exists an open path connecting AA and $B$. Without loss of generality, we suppose $\min|x_{i}-x_{j}|>1$. For $s>0$, let $r_{i}<s$. Define
@@
for some constant $C^{\prime},C^{\prime\prime}\in(0,\infty)$. Since $\lim_{\delta\to 0}\frac{\mathbbm{P}^{\delta}[x_{1}^{\delta}\leftrightarrow x_{2}^{\delta}\leftrightarrow x_{3}^{\delta}\leftrightarrow x_{4}^{\delta}]}{\mathbbm{P}^{\delta}[x_{1}^{\delta}\leftrightarrow x_{2}^{\delta},x_{3}^{\delta}\leftrightarrow x_{4}^{\delta}]}$ tends to 11 as $x_{2}\to x_{3}$, we find $C^{\prime\prime}=C$. Combined with (B.1), we obtain (B.2). ∎